OpenStudy (anonymous):

I need help!!!!!!!!!!!!!!! Factor the expression completely, 6(x+5)^2-5(x+5)+1

OpenStudy (anonymous):

6x(x+5)-(5x(+5))+1=6x(x+5)

OpenStudy (anonymous):

@dopek.i.d.s is incorrect

OpenStudy (anonymous):

I figured... level 48 and 18 at helping people hmmm..:)

OpenStudy (anonymous):

i got to 6x^2 + 55x + 126

OpenStudy (anonymous):

umm....its completely factored but i think you mean you want it simplified...? :)

OpenStudy (anonymous):

x^2 = x*x (x+a)^2 = (x+a)(x+a) c(x+a)= c*x + c*a

OpenStudy (anonymous):

after that i think it can be factored :)

OpenStudy (anonymous):

Well I know the answer it is (2x+9)(3x+14)

OpenStudy (anonymous):

yes so lets see how to do that

OpenStudy (anonymous):

my teacher gave that to me but I do not know how to get it.

OpenStudy (anonymous):

u guys go

OpenStudy (anonymous):

So we need to simplify it, combine like terms, then re-factor it

OpenStudy (anonymous):

i feel stupid now

OpenStudy (anonymous):

6(x+5)^2-5(x+5)+1 6(x^2+10x+25)-5x-25+1 6x^2+55x+126

OpenStudy (anonymous):

This is where it gets difficult to factor.

OpenStudy (anonymous):

x=-b+-_/-b^2-4ac/2a?

OpenStudy (anonymous):

_/-= The square root of.

OpenStudy (anonymous):

Okay:)

OpenStudy (anonymous):

\[6(x+5)^2-5(x+5)+1\]lets simplify this: expand the \((x+5)\) and distribute the \(5\) to the parenthesis\[6(x^2+10x+25)-5x-25+1\]distribute the \(6\)\[6x^2+60x+150-5x-25+1\]combine like terms:\[6x^2+55x+126\]Now , we can factor it...i would use Quadratic equation since i find ti to be the "safest":\[{-55\pm\sqrt{55^2-4(6)(126)}}\over2(6)\]\[{-55\pm\sqrt{3025-3024}}\over12\]\[{-55\pm\sqrt{1}}\over12\]\[{-55\pm1}\over12\]\[{-56\over12}~~~~or~~~~{-54\over12}\]\[x=-{14\over3}~~~~or~~~~x=-{9\over2}\] so that means that its...\[(x+{14\over3})~~~and~~~(x+{9\over2})\]so to make those 'un-fractioned we'll multiply by the denominators:\[\large(3x+14)(2x+9)\] tada! :D

OpenStudy (anonymous):

@Jeans123 hope that helps! :)

OpenStudy (anonymous):

I have been proved wrong. @yummydum that was smart to use the quadratic. He found the solutions for x so he knew what the other terms had to be. Then he got rid of the fractions by multiplying by the denominator

OpenStudy (anonymous):

she* ;)

OpenStudy (anonymous):

Sry just habbit of typing, didn't mean anything by it

OpenStudy (anonymous):

Thank you both of you he and she;)

OpenStudy (anonymous):

no harm :) ...Youre Welcome Austin! :) feel free to ask more anytime you need help (when im there of course :))

OpenStudy (anonymous):

In about five minutes I will:)

OpenStudy (anonymous):

i leave pretty soon so ask now if you can :)